OpenStudy (anonymous):

Solving the following equation will require you to use the quadratic formula. Solve the equation for theta between 0 degrees and 360 degrees  and round your answers to the nearest tenth of a degree. 3sin^2θ-2cosθ-1=0 a. 56.1 degrees, 303.9 degrees b. 56.7 degrees, 303.3 degrees c. 57.5 degrees, 302.5 degrees d. 57.1 degrees, 302.9 degrees e. 56.9 degrees, 303.1 degrees

OpenStudy (anonymous):

Oh, I think I got it. 3sin^2(x)-2cos(x)-1=0 becomes -3cos^2(x)-2cos(x)+2=0

OpenStudy (anonymous):

and then the quadratic formula applied from there. a=-3 b=-2 c=2

OpenStudy (radar):

Using the trig identity\[\cos ^{2} \theta +\sin ^{2} \theta=1\]\ looks like a good approach